Consider a circular coil of radius 6 cm that has 70 loops. Its orientation forms magnetic field lines that are perpendicular to the area of the coil. Suppose the magnetic field changed from 0.20 T to 0.65 T in 5 milliseconds, how much average EMF is induced in the coil?


1
Expert's answer
2022-04-11T12:07:08-0400

The average EMF induced in the coil

"{\cal E}=N\frac{\Delta B}{\Delta t}A"

"{\cal E}=70*\frac{0.65-0.20}{0.005}*3.14*0.06^2
=71\:\rm V"